OpenStudy (anonymous):

when solving this problem do I cross multiply or simply add numerators, my struggle is getting the denominators to be the same... 7/6t + 10/9t add simplify if possible

OpenStudy (anonymous):

to get de denominators the same, you need to multply them by the other denomenator: 7/6 * 9/9 = 63/54 10/9 * 6/6=60/54 7/6t + 10/9t = 63/54 t + 60/54 t =123/54 t

OpenStudy (anonymous):

=2 5/18 t

OpenStudy (anonymous):

thank u, but can u clarify the last sentence is that 2 and 5/18t?

OpenStudy (anonymous):

123/3=41, 54/3=18; I think what she meant to say was that it can be simplified down to 41/18. And, since 41 is prime, it can't go down any further.

OpenStudy (anonymous):

thank u